OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: 2x + y 7 5x – 2y 8 (3, 4) (–6, 3) (–2, –5) (1, –1)

OpenStudy (anonymous):

this is easier than the last one because there are no words

OpenStudy (anonymous):

am i looking for x or y?

OpenStudy (anonymous):

solve for \(y\) in the first equation

OpenStudy (anonymous):

so should i just plug in each one and see which one works?

OpenStudy (anonymous):

\[2x+y=7\] subtract \(2x\) from both sides and get \[y=-2x+7\]

OpenStudy (anonymous):

yes that would work, but we might as well solve

OpenStudy (anonymous):

this is like the question you asked 2 questions ago first solve for (y\) in the first equation, get \(y = -2x+7\) then substitute \(-2x+7\) for \(y\) in the second equation and solve \[2x-2(-2x+7)=8\]

OpenStudy (anonymous):

y=-2x+7 where do we plug that in?

OpenStudy (anonymous):

we plug that in for \(y\) in the second equation, using parentheses

OpenStudy (anonymous):

\[5x-2y=8\] becomes \[5x-2(-2x+7)=8\]

OpenStudy (anonymous):

okay so 10 x^2-4x=8

OpenStudy (anonymous):

i think i did that wron glol

OpenStudy (anonymous):

first step is to multiply everything inside the parentheses by -2 i.e. remove parentheses using distributive property

OpenStudy (anonymous):

then we will combine like terms and then we will start to solve

OpenStudy (anonymous):

you want to try or you want me to write the first step?

OpenStudy (amistre64):

solve for the INequalities; with 4 options, plug and play would be my bet

OpenStudy (anonymous):

5x*-2x= -10x^2-4x=8

OpenStudy (anonymous):

i see the confusion the first term \(5x\) is not being multiplied by the rest

OpenStudy (anonymous):

\[5x-2(-2x+7)=8\] \[5x+4x-14=8\]

OpenStudy (anonymous):

9x-14=8 +14 +14 9x=22 /9 /9 2.4 ? :(

OpenStudy (anonymous):

ok before we go further, that is correct

OpenStudy (anonymous):

so good work on that part i just solved \[2x + y=7\]\[ 5x – 2y=8 \]and got just what you did so there must be a typo in your question the answer is not one of the choices listed

OpenStudy (anonymous):

maybe the question has a typo somewhere in it look carefully and see if the question is correct

OpenStudy (anonymous):

i know what it is. there was supposed to be a \[\ge \] between y and 7 and a \[\le \] between 2y and 8

OpenStudy (anonymous):

\[2x + y\geq7\]\[5x-2u\leq 8\] do what @amistre64 said check which ones work

OpenStudy (amistre64):

yeah, unless you graph it; double checking the options is the simplest route i know if

OpenStudy (amistre64):

 and these things are terrible at graphing :/